# Can Slow-Thinking LLMs Reason Over Time? Empirical Studies in Time Series Forecasting

### v3

@@ 7. Conclusion and Discussions
This work explored whether large language models (LLMs) with slow-thinking capabilities can reason over temporal patterns for time series forecasting. We reformulated forecasting as a conditional reasoning task, encoding historical observations and contextual features into structured prompts to guide future trajectory generation. To realize this paradigm, we proposed TimeReasoner, a prompting-based framework that integrates hybrid inputs—including tabular time series, timestamps, and semantic descriptors—and supports multiple inference-time reasoning strategies (one-shot, decoupled, and roll-out). Experiments across diverse TSF benchmarks demonstrate that slow-thinking LLMs achieve strong zero-shot forecasting performance, while their intermediate reasoning traces reveal coherent patterns of temporal dynamics. This study represents an initial step toward reasoning-centric forecasting systems.
Discussion. Our results demonstrate the potential of slow-thinking approaches for time series forecasting, showing that even a relatively simple framework like TimeReasoner can benefit from the evolving ecosystem of reasoning-oriented LLMs, such as DeepSeek-R1, and is expected to gain further performance improvements as more advanced models (e.g., DeepSeek-R2) become available. However, challenges remain: the reasoning trajectories generated by LLMs still exhibit uncertainty, calling for future work on principled uncertainty quantification; and the existence of erroneous predictions in some cases highlights the need for more robust reasoning and error-correction mechanisms. Addressing these limitations will be essential for realizing reliable and generalizable reasoning-based forecasting systems.
@@ 8. Acknowledgments
This research was supported by grants from the National Natural Science Foundation of China (No. 62502486), the grants of Provincial Natural Science Foundation of Anhui Province (No. 2408085QF193), USTC ResearchFunds of the Double First-Class Initiative, the Fundamental Research Funds for the Central Universities of China.

### v4

@@ 7. Conclusion and Discussions
This work explored whether large language models (LLMs) with slow-thinking capabilities can reason over temporal patterns for time series forecasting. We reformulated forecasting as a conditional reasoning task, encoding historical time series and contextual features into structured prompts to guide trajectory generation. To realize this paradigm, we proposed TimeReasoner, a prompting-based framework that integrates hybrid inputs, including time series, timestamps, and semantic descriptors, further supporting multiple inference-time reasoning strategies (one-shot, decoupled, and roll-out). Experiments across diverse TSF benchmarks demonstrate that slow-thinking LLMs achieve strong zero-shot forecasting performance, while their intermediate reasoning traces reveal coherent patterns of temporal dynamics. This study represents an initial step toward reasoning-centric forecasting systems.
Discussions. Our results demonstrate the potential of slow-thinking approaches for time series forecasting, showing that even a relatively simple framework like TimeReasoner can benefit from the evolving ecosystem of reasoning-oriented LLMs, such as DeepSeek-R1, and is expected to gain further improvements as more advanced models become available. However, challenges remain: the reasoning trajectories generated by LLMs still exhibit uncertainty, calling for future work on principled uncertainty quantification; and the existence of erroneous predictions in some cases highlights the need for more robust reasoning and error-correction mechanisms. Addressing these limitations will be essential for realizing reliable and generalizable reasoning-based forecasting systems.
@@ 8. Acknowledgments
This research was supported by grants from the National Natural Science Foundation of China (No. 62502486, 62337001), the grants of Provincial Natural Science Foundation of Anhui Province (No. 2408085QF193), USTC Research Funds of the Double First-Class Initiative (No. YD2150002501), the Fundamental Research Funds for the Central Universities of China (No. WK2150110032).